12. V 2 O 5 + Cl 2  VOCl 3 + O 2
a. How many moles of V 2 O 5 are required to produce 7.0 moles of VOCl 3 ?
1
Expert's answer
2020-01-23T06:02:17-0500

According to the equation:

2V2O5 + 6Cl2 = 4VOCl3 + 3O2

n(V2O5) = n(VOCl3)/2 = 7/2 = 3.5 moles